If an investor invested $1300 in two mutual funds, one was at 5% and the other at 2%, and a total of $50 was the profit, how much was invested in each fund?
Answer by mananth(16949)

Part I 5.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y
1300
Interest----- 50

Part I 5.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + 1 y= 1300 -------------1
Interest on both investments
5.00% x + 2.00% y= 50
Multiply by 100
5 x + 2 y= 5000.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -6500.00
Add this to (2)
0 x -3 y= -1500
divide by -3
y = 500
Part I 5.00% $ 800
Part II 2.00% $ 500

CHECK
800 --------- 5.00% ------- 40.00
500 ------------- 2.00% ------- 10.00
Total -------------------- 50.00

m.ananth@hotmail.ca